Find the inverse of 35 modulo 11 by using extended Euclidean Algorithm



step by step solution

1
Expert's answer
2021-11-09T11:00:11-0500

"35=3\times11+2...(1)\\ 
11=5\times2+1...(2)\\
2=2\times1+0\\"

Now,

"1=11-5\times2\ \text{From (2)}\\
1= 11-5\times (35-3\times11) \ \text{From (1)}\\
1= 11-5\times35+15\times11\\
1=-5\times35+16\times11\\
1(\text{mod 11})=6\times35+16\times11\\
1(\text{mod 11})=6\times35"

"6\equiv 35^{-1} \text{mod 11}"